In Denver, children bring their old jack-o-lanterns to the top of a tower and compete for accuracy in hitting a target on the ground. Suppose that the tower is 9.0 m high.

A jack-o-lantern is given an initial horizontal speed of 3.4 m/s , what are the direction and magnitude of its velocity 0.80 s later?

A jack-o-lantern is given an initial horizontal speed of 3.4 m/s , what is the direction of its velocity just before it lands?
Answer #1

time of flight

h = 0.5 gt^2

9 = 4.9* t^2

t = 1.355 s

after 0.8 s

vertical velocity

vy = 9.8* 0.8 = 7.84 m/s

net velocity at that instant

v^2 = 3.4^2 + 7.84^2

v = 8.546 m/s

direction

x = arctan ( 7.84/3.4) = 66.55 below horizontal

======

b)

vertical velocity just before hitting ground

vy = 9.8* 1.355 = 13.279 m/s

direction

x = arctan ( 13.279 / 3.4) = 75.638 below horizontal

=====

Comment before rate in case any doubt, will reply for sure.. goodluck